OpenStudy (anonymous):

write 3x-4y=24 in slope intercept form

OpenStudy (jackellyn):

y=mx+b is slope intercept form @Dee_123 So solve for y -4y=24-3x y=-(24/4)-(-3x/4)

OpenStudy (anonymous):

y=3x-24/4

OpenStudy (jackellyn):

@Muskan you have to divide -4 by everything on the right side so the answer would be y=3/4x-6

OpenStudy (anonymous):

k